# UKAN: Unbound Kolmogorov-Arnold Network Accompanied with Accelerated Library

### v2

@@ 3.2 Tasks
To evaluate the performance of UKAN in regression, we conducted three experiments.
@@
    $f(\mathbf{x})=\exp(\frac{1}{15}\sum_{i}^{n}\sin({(\frac{4i}{15}+1)\pi x_{i}}))$ , where $i=0,1,…,15$ and is a high dimensional function compared with functions I and II.
@@
We compare the results of UKAN, KAN and MLP [2, 5, 1] for 2D functions (functions I and II). For function III, we use UKAN, KAN, and MLP [16, 32, 1]. We used a two-layer MLP with 8- and 16-dimensional positional encodings and feature embeddings for 2D and 16D functions, respectively. The first layer of the CG MLP uses SiLU nonlinearity and generated coefficients are scaled by another learnable parameter to improve learning, analogous to the original KAN paper. An Adam optimizer Kingma & Ba (2014) with a learning rate of 0.01 and weight decay of $1e^{-5}$ for 200,000 epochs is used to minimize the MSE loss. The learning rate is decayed exponentially with the rate of $1-1e^{-4}$ and minimum learning rate of $1e^{-4}$. The results are shown in Figure 4, where UKAN and KAN perform much better than MLP, and KAN performs better than UKAN. In theory, KAN and UKAN have the same learning capacity, but the MLP component of UKAN might slightly hurt generalization and performance compared to KAN. One might switch to a KAN-based CG model instead of an MLP-based CG model to resolve the generalization issues, but this comes with large computational costs. In terms of a fixed FLOPs budget, our study provides a more robust and fair comparison in terms of KAN and MLP accuracy, even though this was not a target of our study. We additionally note that although the compute cost of a KAN is greater than that of an MLP by a factor of $K$, KAN convergence often compensates for this factor.
@@
Figure 4: The regression task results. a. RMSE over training epochs of function I with KAN, UKAN, and MLP [2, 5, 1]. b. RMSE over training epochs of function II with KAN, UKAN, and MLP [2, 5, 1]. c. RMSE over training epochs of function III with KAN, UKAN, and MLP [16, 32, 1].
@@
KANs promise better generalization compared to MLPs for regression tasks, similar to equivariant models allowing for the exploitation of symmetries for improved generalization. In particular, E(n)-Equivariant Graph Neural Networks (EGNNs) are equivariant with respect to the translations, rotations, and permutations Satorras et al. (2022). Here, we explore how combining equivariance with KAN leads to improved performance in the study of n-body systems as described in the EGNN paper Satorras et al. (2022). To evaluate this, we replace the final scalar predicting MLPs in EGNN with UKAN and KAN layers. Specifically, the scalar outputs of $\phi_{x}$ and $\phi_{v}$ in Equation 4 are predicted with UKAN and KAN.
$\mathbf{v}_{i}^{l+1}$ $=\phi_{v}(\mathbf{h}_{i}^{l})\mathbf{v}_{i}^{init}+C\sum_{j\neq i}(\mathbf{x}_{i}^{l}-\mathbf{x}_{j}^{l})\phi_{x}(\mathbf{m}_{ij}^{l+1})$ (4)
$\mathbf{x}_{i}^{l+1}$ $=\mathbf{x}_{i}^{l}+\mathbf{v}_{i}^{l+1}$
@@
We keep the rest of parameters and datasets identical to the original paper and their code on Github. We also train the SE(3) Transformer model as another reference point. The results are shown in Table 2, where we observe that UKAN and KAN improve the accuracy compared to the original architecture, and the improvement of UKAN is better than the KAN model.
@@
We explore the effectiveness of UKAN and KAN in physics-informed neural networks Karniadakis et al. (2021) to solve the logistic growth model, which is used to model population dynamics in biological and ecological systems. For this experiment, we use both UKAN and KAN [1, 5, 1] without MLP component to solve the differential equation below,
$\frac{df}{dt}=Rf(t)(1-f(t))$ (5)
where $R$ is the growth rate set to 1.0 and the function $f(t)$ represents the growth rate of the population over time (t). We impose boundary condition of $f(0)=0.5$ to uniquely specify the solution and compare the results with the analytical solution of $f(t)=\frac{1}{1+\exp{(-t)}}$. We use domain of $[-5,5]$ to sample data and Adam optimizer with the learning of rate of $1e^{-3}$ and weight decay of $1e^{-5}$ and follow the standard procedure for PINN minimization, i.e. minimizing MSE of the differential equation residual over colocation points and boundary conditions. The results are shown in Figure 5, for and KAN [1, 5, 1] without the MLP component. UKAN and KAN acheive MSE of $1e^{-5}$ and $1e^{-6}$, respectively on the sample dataset, indicating both models are applicable to physics-informed neural networks scenarios.
Figure 5: KAN and UKAN used in PINNs. Solving logistic growth model with both KAN and UKAN [1, 5, 1] over domain of -5 and 5. Both models match the analytical solution with good approximation.
We devised two experiments to evaluate UKAN’s performance in classification tasks. In particular, we trained UKAN and KAN over Moon- and MNIST-datasets. We report the accuracy of training and validation of both models on the moon-dataset in Table 3, where we observe close to 100% accuracy for Moon-dataset with 2D inputs with slight superior accuracy of UKAN over KAN in this setup. Both UKAN and KAN [2, 4, 2] are trained using SGD optimizer with learning rate of 0.01 for 10000 epochs, and results are averaged over 3 different initializations.
Table 3: Moon dataset classification accuracy.
Model Training Validation
KAN $98.46\pm 1.3$ $98.53\pm 0.4$
UKAN $\mathbf{100.0\pm 0.}$ $\mathbf{99.83\pm 0.17}$
The final classification task was performed on the MNIST dataset, where we trained both the UKAN and KAN models with configurations [784, 32, 10] and a degree-3 B-spline. Both models were optimized using the Adam optimizer combined with an Exponential scheduler, having a learning rate of $2\times 10^{-4}$ and a decay rate of $1-10^{-4}$. The KAN network incorporated 51 grid points across the interval $[-10,10]$, whereas UKAN utilized a grid delta of 3.0 and a 48-dimensional positional encoding. Notably, both models employed only the B-spline component without any MLP components.
Figure 6 illustrates the learning curves for both the training and validation datasets. Training was halted upon detection of overfitting in the training dataset. Furthermore, three rounds of independent training with different initializations were conducted to compare the performance of UKAN and KAN. The results, as presented in Table 4, indicate that UKAN outperforms KAN on the validation dataset while slightly underperforming on the training dataset.
Table 4: MNIST dataset classification accuracy.
Model Training Validation
KAN $\mathbf{98.93\pm 0.78}$ $95.35\pm 0.04$
UKAN $98.40\pm 0.24$ $\mathbf{96.29\pm 0.08}$
Figure 6: Classification task over MNIST dataset. Accuracy of training and validation datasets for UKAN and KAN [784, 32, 10] as training progresses. UKAN uses a 48-dimensional positional embedding and feature index embedding.
Finally, we evaluated the performance of three different architectures for Denoising Diffusion Probabilistic Models (DDPM) Ho et al. (2020) on a synthetic 2D circle dataset with added noise. The first architecture is only composed of MLPs, while other architectures use KAN and UKAN in the input of temporal layers and output layer (see Appendix A for full architecture). We used Adam optimizer with a learning rate of $5e^{-5}$ for 500 epochs with a batch size of 800. Our results demonstrated that both KAN and UKAN significantly outperformed MLP in terms of the Wasserstein distance shown in Table 5 and sample quality as shown in Figure 7. Data samples from original distribution and generated from DDPM with KAN, UKAN, and MLP architectures indicates superior performance of KAN and UKAN compared to MLP and slightly superior performance of UKAN over KAN. This result indicates possible applications of KAN and UKAN in generative tasks, where MLP alone might fail to learn underlying data distribution especially in sample quality as we observed loss values of MLP, KAN and UKAN were very small.
Table 5: DDPM with KAN, UKAN, and MLP
Model Wasserstein distance
KAN $0.693$
UKAN $\mathbf{0.655}$
MLP $1.058$
Figure 7: DDPM with KAN, UKAN, and MLP.
4 conclusion
In this work, we presented the Unbound Kolmogorov-Arnold Network (UKAN) which unifies multi-layer perceptron networks (MLPs) with KANs along with an efficient GPU implementation of the underlying components of KANs. GPU acceleration decouples the computational cost and memory fingerprint of KAN from the grid size to the order of B-spline function by using local matrix representations of B-spline functions. In addition, our proposed UKAN architecture allows using KANs without any fixed grid range limitation by generating coefficients from a coefficient-generator MLP. We evaluated our UKAN model on regression, classification and generative tasks and compared learning efficiency and accuracy. We are excited about the potential of UKAN and its variants, as well as the applications of our accompanying GPU optimized library which alleviates memory and compute issues in existing implementations. We believe this work will enable application of accelerated KAN and UKAN in areas such as molecular property predictions, protein docking, large language models, and computer vision.
Appendix A Appendix

### v4

@@ 3.2 Tasks
3.2.1 Regression
To evaluate the accuracy of our UKAN and KAN in regression, we conducted three experiments.
@@ 3.2.1 Regression
    $f(\mathbf{x})=\exp(\frac{1}{15}\sum_{i}^{n}\sin({(\frac{4i}{15}+1)\pi x_{i}}))$ , where i=0,1,,15i=0,1,…,15 and is a high dimensional function compared with functions I and II.
@@
We compare the results of UKAN, KAN and MLP $(2\!\to\!5\to\!1)$ for 2D functions (functions I and II). For function III, we use UKAN, KAN, and MLP $(16\!\to\!32\to\!1)$. For the CG model, we used a two-layer MLP with 8- and 16-dimensional positional encodings and feature embeddings for 2D and 16D functions, respectively. The first layer of the CG MLP uses SiLU nonlinearity and generated coefficients are scaled by another learnable parameter to improve learning, analogous to the original KAN paper. An Adam optimizer Kingma & Ba (2014) with a learning rate of 0.01 and weight decay of $1e^{-5}$ for 200,000 epochs is used to minimize the MSE loss. The learning rate is decayed exponentially with the rate of $1-1e^{-4}$ and minimum learning rate of $1e^{-4}$. The results are shown in Figure 3, where UKAN and KAN perform much better than MLP, and KAN performs better than UKAN. In theory, KAN and UKAN have the same learning capacity, but the MLP component of UKAN might slightly hurt generalization and performance compared to KAN. We also note that although the compute cost of a KAN is greater than that of an MLP by a factor of $K$, KAN convergence often compensates for this factor. We also performed a regression task on the n-body problem based on Satorras et al. (2022), the results are available in Appendix A.3.
@@
Figure 3: Regression task results. (a) RMSE vs. training epochs for Function I using KAN, UKAN, and $\mathrm{MLP}(2\!\to\!5\!\to\!1)$. (b) RMSE vs. training epochs for Function II using KAN, UKAN, and $\mathrm{MLP}(2\!\to\!5\!\to\!1)$. (c) RMSE vs. training epochs for Function III using KAN, UKAN, and $\mathrm{MLP}(16\!\to\!32\!\to\!1)$.
3.2.2 Classification
We devised two experiments to evaluate UKAN’s performance in classification tasks. In particular, we trained UKAN and KAN over Moon- and MNIST-datasets. We report the accuracy of training and validation of both models on the moon-dataset in Table 3, where we observe close to 100% accuracy for Moon-dataset with 2D inputs with slight superior accuracy of UKAN over KAN in this setup. Both UKAN and KAN $(2\!\to\!4\to\!2)$ are trained using SGD optimizer with learning rate of 0.01 for 10000 epochs, and results are averaged over 3 different initializations.
Table 3: Moon dataset classification accuracy.
Model Training Validation
KAN $98.46\pm 1.3$ $98.53\pm 0.4$
UKAN $\mathbf{100.0\pm 0.}$ $\mathbf{99.83\pm 0.17}$
The final classification task was performed on the MNIST dataset, where we trained both the UKAN and KAN models with configurations $(784\!\to\!32\to\!10)$ and a degree-3 B-spline. Both models were optimized using the Adam optimizer combined with an Exponential scheduler, having a learning rate of $2\times 10^{-4}$ and a decay rate of $1-10^{-4}$. The KAN network incorporated 51 grid points across the interval [10,10][-10,10], whereas UKAN utilized a grid delta of 3.0 and a 48-dimensional positional encoding. Notably, both models employed only the B-spline component without any MLP components. Training was halted upon detection of overfitting in the training dataset. Furthermore, three rounds of independent training with different initializations were conducted to compare the performance of UKAN and KAN. The results, as presented in Table 4, indicate that UKAN outperforms KAN on the validation dataset while slightly underperforming on the training dataset.
Table 4: MNIST dataset classification accuracy.
Model Training Validation
KAN $\mathbf{98.93\pm 0.78}$ $95.35\pm 0.04$
UKAN $98.40\pm 0.24$ $\mathbf{96.29\pm 0.08}$
3.2.3 Approximation
We explore the effectiveness of UKAN and KAN in physics-informed neural networks Karniadakis et al. (2021) to solve the logistic growth model, which is used to model population dynamics in biological and ecological systems. For this experiment, we use both UKAN and KAN ($1\rightarrow 5\rightarrow 1$) without MLP component to solve the differential equation below,
$\frac{df}{dt}=Rf(t)(1-f(t))$ (15)
where $R$ is the growth rate set to 1.0 and the function $f(t)$ represents the growth rate of the population over time (t). We impose boundary condition of $f(0)=0.5$ to uniquely specify the solution and compare the results with the analytical solution of $f(t)=\frac{1}{1+\exp{(-t)}}$. We use domain of [5,5][-5,5] to sample data and Adam optimizer with the learning of rate of $1e^{-3}$ and weight decay of $1e^{-5}$ and follow the standard procedure for PINN minimization, i.e. minimizing MSE of the differential equation residual over collocation points and boundary conditions. The results are shown in Figure 4, for and KAN $(1\!\to\!5\to\!1)$ without the MLP component. UKAN and KAN achieve MSE of $1e^{-5}$ and $1e^{-6}$, respectively on the sample dataset, indicating both models are applicable to physics-informed neural networks scenarios.
Figure 4: KAN and UKAN used in PINNs. Solving logistic growth model with both KAN and UKAN $(1\!\to\!5\to\!1)$ over domain of -5 and 5.
3.2.4 Generation
As an example in generative model learning, we evaluated the performance of three different architectures for Denoising Diffusion Probabilistic Models (DDPM) Ho et al. (2020) on a synthetic 2D circle dataset with added noise. The first architecture is only composed of MLPs, while other architectures use KAN and UKAN in the input of temporal layers and output layer (see Appendix A.4 for full architecture). We used Adam optimizer with a learning rate of $5e^{-5}$ for 500 epochs with a batch size of 800. Our results demonstrated that both KAN and UKAN significantly outperformed MLP in terms of the Wasserstein distance shown in Table 5 and sample quality as shown in Figure 5. Data samples from original distribution and generated from DDPM with KAN, UKAN, and MLP architectures indicates superior performance of KAN and UKAN compared to MLP and slightly superior performance of UKAN over KAN. This result indicates possible applications of KAN and UKAN in generative tasks, where MLP alone might fail to learn underlying data distribution especially in sample quality as we observed loss values of MLP, KAN and UKAN were very small.
Table 5: DDPM with KAN, UKAN, and MLP
Model Wasserstein distance
KAN $0.693$
UKAN $\mathbf{0.655}$
MLP $1.058$
Figure 5: DDPM with KAN, UKAN, and MLP.
3.2.5 Real-world Application: Drug Discovery
Accurate in silico prediction of molecular properties is key to accelerating drug discovery by early identification of compounds with favorable ADME (Absorption, Distribution, Metabolism, and Excretion) profiles. Ferreira & Andricopulo (2019); Beckers et al. (2023); Seal et al. (2025) Machine learning methods have shown strong performance in predicting these properties from chemical structure or computed features. The usage of KANs in predicting molecular properties was originally introduced by Li et al. (2025). Due to implementation discrepancies and known issues within the MoleculeNet datasets,Walters we reimplemented the original model using the warpKAN package and incorporated the necessary corrections. A comprehensive analysis of these results is presented in the Appendix C.
Here, we further demonstrate the ability of UKANs to predict ADME properties, and used a fixed molecular representation to isolate the impact of the KAN and UKAN addition. We choose Morgan fingerprints and RDKit 2D descriptors as our two computed molecular features. In order to overcome the limitations of the MoleculeNet dataset, we use four datasets released by Fang et al. (2023). These models were trained for 200 epochs using Adam optimizer with a learning rate of 1e-4. As shown in Table 6, UKANs show superior or comparable performance with respect to KANs for predicting molecular properties from computed features using mean absolute error metric. UKANs outperform KANs on the permeability and solubility dataset using Morgan fingerprint features while their performance is statistically equivalent (using T-test for means) on the other two datasets. Using RDKit 2D descriptors, a set of 200 molecular descriptors describing the 2D structure of the molecule, UKAN shows better performance than KAN on the Microsomal stability (rat) dataset. These results show that UKANs are superior to KANs in accurately predicting molecular properties relevant to the pharmaceutical industry. Details about model architecture and data preprocessing can be found in the Appendix A.5 and B.1 respectively.
Table 6: Performance of KAN and UKAN on Biogen ADME datasets using Morgan fingerprint (labeled Morgan) and RDKit 2D descriptors (labeled RDKit-2D) as molecule featurizers. Performance is assessed using mean absolute error on test split as the metric. ($\downarrow$ is better)
Descriptor Microsomal Microsomal Permeability Solubility
stability (human) stability(rat)
KAN Morgan $0.479\pm 0.010$ $0.559\pm 0.012$ $0.424\pm 0.005$ $0.476\pm 0.006$
UKAN Morgan $0.476\pm 0.003$ $0.551\pm 0.020$ $\mathbf{0.403\pm 0.006}$ $\mathbf{0.435\pm 0.009}$
KAN RDKit-2D $0.373\pm 0.003$ $0.492\pm 0.010$ $0.349\pm 0.005$ $0.386\pm 0.002$
UKAN RDKit-2D $0.368\pm 0.008$ $\mathbf{0.467\pm 0.006}$ $0.348\pm 0.003$ $0.397\pm 0.013$
4 conclusion
In this work, we presented the Unbound Kolmogorov-Arnold Network (UKAN), which unifies multilayer perceptron networks (MLPs) with KANs along with an efficient GPU implementation of the underlying components of KANs. GPU acceleration decouples the computational cost and memory fingerprint of KANs from the grid size by using local matrix representations of B-spline functions. In addition, our proposed UKAN architecture allows the use of KANs without any fixed grid range limitation by generating coefficients from a coefficient-generator MLP. We evaluated UKAN model for regression, classification, and generative tasks. Our accompanying GPU library alleviates the core bottlenecks that have limited the practical scale of spline-based networks and serves as a reusable building block for future models. We expect UKAN and its variants to enable accelerated large-scale learning in domains such as molecular property prediction, protein docking, language, and vision, and we see promising directions in multi-GPU training, adaptive knot policies, and a deeper theory of approximation on unbounded domains.
Appendix A Appendix
You may include other additional sections here.
A.1 Cubic B-Spline Basis Matrix Representation
For a B-spline of order 3, the basis matrix representation can be written as,
$\operatorname{spline}(u)=\begin{bmatrix}1&u&u^{2}&u^{3}\end{bmatrix}\frac{1}{6}\begin{bmatrix}1&4&1&0\\ -3&0&3&0\\ 3&-6&3&0\\ -1&3&-3&1\end{bmatrix}\begin{bmatrix}c_{0}\\ c_{1}\\ c_{2}\\ c_{3}\end{bmatrix}$.
A.2 UKAN mapping to KAN in Inference
Proposition (UKAN $\rightarrow$ KAN at inference).
Fix B–spline degree $k$ and uniform grid spacing $h$, and set $K:=k+1$. Consider a UKAN whose coefficient–generator (CG) produces a group vector $G_{g}\in\mathbb{R}^{K}$ for each group index $g\in\mathbb{Z}$. Evaluation uses the same local stencil as KAN,
$\text{spline}(x)=U(u)\,M_{k}\,(\cdot),\qquad U(u)=(1,u,\ldots,u^{k}),\ \ u=\tfrac{x-t_{i}}{h}\in[0,1),$
with $t_{i}=i\,h$ the knot at cell index $i$. Let $\mathcal{X}\subset\mathbb{R}$ be the (finite) set of inputs on which inference is performed; more generally, let $\mathcal{I}$ be any compact interval containing $\mathcal{X}$. Then there exists a KAN with coefficients $\{C_{i}\}$ defined only for cells intersecting $\mathcal{I}$ such that the KAN exactly matches the UKAN on $\mathcal{X}$ (indeed, on all of $\mathcal{I}$).
Proof. For each cell index $i$ whose cell $t_{i},t_{i+1})$ intersects $\mathcal{I}$, write $i=gK+r$ with
$g=\Big\lfloor\tfrac{i}{K}\Big\rfloor,\qquad r=i\bmod K\in\{0,\ldots,K-1\}.$
Query the CG once (in inference mode; parameters fixed) to obtain the adjacent group vectors $G_{g-1},G_{g}\in\mathbb{R}^{K}$, and materialize the KAN’s local coefficients by the sliding window
$C_{i}\;:=\;\mathrm{slice}\!\left(\,[G_{g-1};\,G_{g}],\;r:\,r+K\,\right)\in\mathbb{R}^{K},$
where $\cdot\,$$\cdot\,$] denotes concatenation and $\mathrm{slice}(\cdot,a{:}b)$ extracts entries $a,\ldots,b-1$. By construction, for any $x$ in cell $i$ the UKAN evaluation applies the same linear stencil $U(u)M_{k}$ to the same length-$K$ window drawn from $G_{g-1};G_{g}]$. Hence
$U(u)M_{k}\,C_{i}\;=\;U(u)M_{k}\,\mathrm{slice}\!\left(\,[G_{g-1};G_{g}],\,r{:}r+K\right),$
which is exactly the UKAN output in that cell. Because $\mathcal{I}$ intersects finitely many cells, the above defines finitely many $C_{i}$ and therefore a finite–parameter KAN that agrees with the UKAN on all $x\in\mathcal{I}$, in particular on the finite inference set $\mathcal{X}$.
A.3 N-body Problem
@@
KANs promise better generalization compared to MLPs for regression tasks, similar to equivariant models allowing for the exploitation of symmetries for improved generalization. In particular, E(n)-Equivariant Graph Neural Networks (EGNNs) are equivariant with respect to the translations, rotations, and permutations Satorras et al. (2022). Here, we explore how combining equivariance with KAN leads to improved performance in the study of n-body systems as described in the EGNN paper Satorras et al. (2022). To evaluate this, we replace the final scalar predicting MLPs in EGNN with UKAN and KAN layers. Specifically, the scalar outputs of $\phi_{x}$ and $\phi_{v}$ in equations below are predicted with UKAN and KAN.
$\mathbf{v}_{i}^{\ell+1}$ $=\phi_{v}(\mathbf{h}_{i}^{\ell})\,\mathbf{v}_{i}^{\mathrm{init}}+C\sum_{j\neq i}(\mathbf{x}_{i}^{\ell}-\mathbf{x}_{j}^{\ell})\,\phi_{x}(\mathbf{m}_{ij}^{\ell+1}),\qquad\mathbf{x}_{i}^{\ell+1}=\mathbf{x}_{i}^{\ell}+\mathbf{v}_{i}^{\ell+1}.$
@@
Where $\mathbf{x}_{i}^{(\ell)}$ and $\mathbf{v}_{i}^{(\ell)}$ are the position and velocity of $i^{th}$ particle in the $l^{th}$ layer of EGNN. We keep the rest of parameters and datasets identical to the original paper and their code on Github. We also train the SE(3) Transformer model as another reference point. The results are shown in Table 7, where we observe that UKAN and KAN improve the accuracy compared to the original architecture, and the improvement of UKAN is better than the KAN model.
@@
A.4 Architecture Details of DDPM
@@
A.5 Architecture details of molecular property prediction
The model architecture consists of 3 layers of KAN/UKAN with hidden dimension equal to 2X the dimension of the input computed molecular features. The Morgan fingerprint was computed using RDKit packageLandrum et al. (2016) with a radius of 2 and a bit length of 1024, yielding a molecular features of dimension 1024. RDKit 2D descriptors were computed using RDKit package and normalized using descriptastorus packagedes resulting in a vector of size 200. Both KAN and UKANs were used with order 3 B-spline functions and a grid delta of $1.0$.
A.6 Architecture Details of GCN
KAN/UKAN/FKAN-GCN
We build three GCN variants by replacing MLP block (node embedding, message passing, and readouts) with a KAN, UKAN, or FKAN layer, respectively.
Node embedding. Given initial node features $f_{v}\in\mathbb{R}^{d_{in}}$ and neighbor set $\mathcal{N}(v)$, we form an embedding by concatenating $f_{v}$ with a degree-normalized neighbor average and passing it through a basis layer $\Phi_{E}(\cdot)$ implemented by KAN/UKAN/FKAN:
$h_{v}^{(0)}=\Phi_{E}\left(\big[f_{v};\tfrac{1}{|\mathcal{N}(v)|}\sum_{u\in\mathcal{N}(v)}f_{u}\big]\right)$
Message passing.
At layer $\ell$, the neighbor is first transformed by the basis layer $\Phi_{M}^{\ell}(\cdot)$ and aggregated over the neighbor with summation.
$m_{uv}^{(\ell)}\;=\;\Phi_{M}^{(\ell)}\!\big(h_{u}^{(\ell)}\big),\qquad h_{v}^{(\ell+1)}\;=\;\sum_{u\in\mathcal{N}(v)}m_{vu}^{(\ell+1)}.$
Readout. After $L$ layers (four layers in our experiments), node features are pooled with avg (any permutation-invariant $P\in\{\textsc{avg},\textsc{sum},\textsc{max}\}$ is supported) and mapped to outputs by a KAN/UKAN/FKAN readout:
$\bar{h}\;=\;P\big(\{h_{v}^{(L)}:v\in V\}\big),\qquad\hat{y}\;=\;\Phi_{R}(\bar{h}).$
Here, each $\Phi_{\{\!E,M,R\!\}}(\cdot)$ denotes the same structural layer instantiated with a different basis: KAN uses fixed, bounded B-spline grids; UKAN uses a coefficient-generator to produce local B-spline coefficients on an unbounded symmetric grid; FKAN replaces the spline basis with a Fourier basis. All variants are drop-in compatible.
Appendix B Data prepreprocessing
B.1 Molecular property prediction
The molecular property prediction dataset was curated using standard practices in cheminformatics like removal of invalid SMILES, molecule standardization, removal salts, charge neutralization, and mean aggregation of any duplicate labels. This curation practice roughly following the guidelines outlined in Fourches et al. (2010). The molecules were split into train, validation, and test set using Bemis-Murcko scaffold split for each of the tasks in the dataset. In order to make statistically significant comparisons, only tasks containing a more than 500 labels were used. This resulted in selecting 4 out of the 6 tasks presented in the original paper.Fang et al. (2023)
Appendix C KAN/UKAN/FKAN-GCN results on MoleculeNet
Li et al. (2025) introduced a graph neural network with multi-layer perceptron layers replaced with KANs. The authors demonstrate that such an architectures shows strong performance on the MoleculeNet datasets. However, we identified two issues with the KAN-GCN implementation by Li et al. : (1) Best model selection is performed on test loss instead of validation loss, which is a deviation from the best practices in machine learning. (2) MoleculeNet consists of multi-task datasets with missing labels. The missing labels were assigned a value of $0.0$ instead of being treated as a missing label. Train and test metrics were computed based on this artificial label. We have fixed these issues in our reimplementation using warpKAN library.
To assess scalability of our implementation, we evaluate KAN, UKAN, and Fourier-based KAN (FKAN) graph convolutional neural network (GCN) on seven MoleculeNet datasets Wu et al. (2018). The total number of components in the dataset is over 148,000 molecules, where each molecule is composed of several heavy atoms, e.g., BACE and Tox21 have on average 65 and 36 atoms. We modify a standard GCN Wang et al. (2023) by replacing its MLP blocks with KAN or UKAN layers (see Appendix A.6). To the best of our knowledge, our training is one of the large-scale training that incorporates KAN in all parts of GCN. Note while works like Zhang & Zhang (2024); Bresson et al. (2025) claim KAN usage it is usually limited to initial node embedding or readouts, as without efficient implementations like ours it is infeasible to investigate KAN in large scale. We train three models, where we divide data into training, validation, and testing datasets and select the best model based on the validation loss. We report the area under the curve (AUC) for models in Table 8. Each model is trained for 500 epochs using Adam optimizer with learning rate of 2e-4 and a step scheduler with a decay rate of 0.9 every 20 steps. Unlike values reported in Li et al. (2025), we see the accuracy of B-Spline- and Fourier-based models are very close to each other, when model selection is done correctly.
Table 8: Performance of KAN, UKAN, and FKAN on MoleculeNet datasets ($\uparrow$ is better).
Dataset BBBP BACE ClinTox Tox21 SIDER HIV MUV
Tasks 1 1 2 27 12 1 17
KAN $65.6(2.2)$ $80.0(2.4)$ $96.5(0.2)$ $79.4(0.7)$ $83.5(0.2)$ $76.5(1.3)$ $71.8(5.7)$
UKAN $64.5(2.2)$ $74.9(6.3)$ $94.7(0.9)$ $77.4(0.5)$ $83.7(0.2)$ $75.2(1.0)$ $69.6(2.1)$
FKAN $67.0(2.0)$ $80.5(3.0)$ $96.5(0.1)$ $79.4(0.8)$ $82.2(0.1)$ $74.1(2.2)$ $76.5(1.2)$
Appendix D LLM Usage
We used large-language models (LLMs) as general-purpose assistive tools during the preparation of this paper. Specifically:
  • Writing support: LLMs were used to improve grammar, clarity, and text flow. All technical content, methodology, experimental design, and conclusions were conceived and verified by the authors.
  • Editing and formatting: LLMs assisted in rephrasing sentences for readability, generating LaTeX table and figure formatting, and ensuring consistency in notation.
  • Brainstorming: LLMs were used to explore alternative phrasings, organizational structures, and to verify the completeness of the literature-related sections.
No LLM-generated text or ideas were included without careful review and verification by the authors. The models did not contribute to scientific novelty, research ideation, or experimental results. The authors assume full responsibility for all content in this manuscript.